# A Novel Pathology Foundation Model by Mayo Clinic, Charité, and Aignostics

### v1

@@ Abstract
Recent advances in digital pathology have demonstrated the effectiveness of foundation models across diverse applications. In this report, we present a novel vision foundation model based on the RudolfV approach. Our model was trained on a dataset comprising 1.21.21.2 million histopathology whole slide images, collected from two medical institutions: Mayo Clinic and Charité - Universtätsmedizin Berlin. Comprehensive evaluations show that our model achieves state-of-the-art performance across twenty-one public benchmark datasets, even though it is neither the largest model by parameter count nor by training dataset size.
@@ 1 Introduction
However, despite artificial intelligence (AI) having led to promising proof-of-concepts and applications (e.g., klauschen2024toward ; diao2021human ; raciti2023clinical ; binder2021morphological ; keyl2022patient ), generalization, application variety, and robustness remain challenging and have hindered the broad translation of AI applications into clinical routine diagnostics. Here, the limited adoption of digitization in clinical practice results in a scarcity of training data, particularly for infrequent and rare diseases dippel2024ai . Furthermore, generating sufficient labeled data representing the full spectrum of human disease, biological, and technical variability inherent to morphology, tissue processing, staining, and slide scanners has proven logistically and financially challenging.
@@
In this report, we utilized a corpus of 1.2 million histopathology whole slide images derived from more than 490,000 cases from Mayo Clinic and Charité - Universitätsmedizin Berlin to train a ViT-H/14 pathology foundation model using the training paradigm from RudolfV dippel_rudolfv_2024 . Our model incorporates a broad diversity of diseases, staining types, and scanners, and utilizes multiple image magnifications during training. We provide an assessment of its performance compared to other leading models available for testing using twenty-one benchmark datasets assessing a variety of downstream pathology tasks. An overview of model characteristics and results can be found in Figure 1.
Figure 1: Overview of average performance, training dataset size, and model size of different contenders. The average performance is detailed in Table 1. H-Optimus-0 hoptimus0 and Prov-GigaPath xu_whole-slide_2024 are the models with the most parameters and Virchow2 zimmermann_virchow_2024 is the model trained on most slides. Our model exhibits the best average performance and an intermediate model and training dataset size.
@@ 2 Related Work
Multiple previous works on histopathology foundation models have developed a variety of models with increasingly large parameter counts and on increasingly expansive pathology datasets curated from public and private sources dippel_rudolfv_2024 ; zimmermann_virchow_2024 ; filiot_phikon-v2_2024 ; xu_whole-slide_2024 ; hoptimus0 ; chen_towards_2024 ; ding2024multimodalslidefoundationmodel ; lenz2024unsupervisedfoundationmodelagnosticslidelevel ; juyal_pluto_2024 ; lu2024avisionlanguage ; qiu2024pixelsgigapixelsbridginglocal . Pathology foundation models can be separated into tile- and slide-based models. Tile-based models, which represent the majority of the models, generate embeddings from fixed-sized image tiles derived from gigapixel whole slide images (WSIs). Tile-based models have typically been trained via self-supervised-learning dippel_rudolfv_2024 ; zimmermann_virchow_2024 ; filiot_phikon-v2_2024 ; xu_whole-slide_2024 ; hoptimus0 ; chen_towards_2024 ; ai_towards_2024 ; nechaev_hibou_2024 ; wang_transformer-based_2022 ; kang2022benchmarking ; lu2024avisionlanguage and most studies dippel_rudolfv_2024 ; zimmermann_virchow_2024 ; filiot_phikon-v2_2024 ; xu_whole-slide_2024 ; hoptimus0 ; chen_towards_2024 ; ai_towards_2024 ; nechaev_hibou_2024 are based on the DINOv2 framework oquab2023dinov2 . We include the leading and available tile-based models in our study, i.e. Virchow2 (632 million parameters; 3.1 million slides; zimmermann_virchow_2024 ), H-Optimus (1.1 billion parameters; 500k slides; hoptimus0 ), RudolfV (300 million parameters; 134k slides; dippel_rudolfv_2024 ), UNI (300 million parameters; 100k slides; chen_towards_2024 ), and Phikon 2 (300 million parameters; 58k slides; filiot_phikon-v2_2024 ).
Slide-based foundation models seek to create global representations of whole slide images, which requires encoding features that may span individual tiles or otherwise represent global phenomena not readily evident on single tiles. To address this, most current slide-based models operate in two steps ding2024multimodalslidefoundationmodel ; lenz2024unsupervisedfoundationmodelagnosticslidelevel ; wang_pathology_2024 ; xu_whole-slide_2024 ; shaikovski_prism_2024 . First, by encoding individual tiles and secondly by aggregating the resulting tile representations into a slide-level representation. For such models, the tile-encoder is a key performance gateway. Accordingly, this study compares the tile-based encoder of Prov-GigaPath (1.1 billion parameters; 171k slides; xu_whole-slide_2024 ).
To date, many foundation models in pathology have been trained exclusively on hematoxylin and eosin (H&E) stained slides with tiles extracted at a single magnification level. While H&E staining represents the bulk of routine pathology, alternative histochemical and immunohistochemical (IHC) stains are the basis for many biomarker evaluations and accordingly have been increasingly represented in recent studies zimmermann_virchow_2024 ; dippel_rudolfv_2024 ; xu_whole-slide_2024 . Similarly, data from multiple image magnifications have been included in recent work zimmermann_virchow_2024 ; ai_towards_2024 . The presented model incorporates these features, and uses data obtained from H&E, IHC, and special stains, as well as multiple magnifications.
@@ 3.1 Dataset and Preprocessing
A curated set of 1.2 million de-identified WSIs, derived from 490k pathology cases, from the digital archives of Mayo Clinic and Charité - Universitätsmedizin Berlin, was utilized to generate 3.4 billion image tiles for training. Tiles were extracted at multiple resolutions, namely 0.25, 0.5, 1.0, and 2.0 microns per pixel, corresponding to objective microscopic magnifications of 40×40$\times$40 ×, 20×20$\times$20 ×, 10×10$\times$10 ×, and 5×5$\times$5 ×, respectively. An overview of relevant dataset statistics are given in Figure 2. Slides and tiles were sampled for training using the sampling algorithm of RudolfV dippel_rudolfv_2024 .
@@ 3.2 Model Framework and Compute Environment
A sampled dataset of ca. 520m tiles was used to train a ViT-H/14 (632 million parameters) model dosovitskiy2020vit using an adapted RudolfV dippel_rudolfv_2024 approach, which is based on the DinoV2 framework oquab2023dinov2 . Model training was performed with Nvidia H100 GPUs within the Mayo Clinic Platform environment [^1] https://www.mayoclinicplatform.org/.
@@ 3.3 Evaluation Protocols
We evaluated model performance using linear probing protocols as established in the literature dippel_rudolfv_2024 ; hoptimus0 ; zimmermann_virchow_2024 , with all models evaluated on extracted embeddings from frozen backbones. We use both public benchmarks as well as public evaluation frameworks where available, to foster reproducibility and comparability. Results were computed for 5 seeds over data split, shuffling, and “probing” model initialization per foundation model and task, if not stated otherwise in a specific task description (see Appendix A.1). We report the mean performance and standard errors over the seeds. Dataset splits are described in detail in the respective task descriptions in Appendix A.1. No augmentations were applied when extracting embeddings. As all models use Vision Transformer architectures dosovitskiy2020vit , we always evaluate both the CLS token and the CLS+Mean [^2] CLS+Mean being the concatenation of the CLS token and the mean over all “mini-patch” tokens token embeddings for every model and report the better (maximum) performance of the two in Table 1. This accounts for potential systematic differences between information encoded in different tokens between different models, as also done in previous works (see e.g. zimmermann_virchow_2024 ). Results of using the CLS and CLS+Mean token only, are reported in Table 4 and Table 5, respectively.
@@
We use eva kaiko.ai2024eva , an open-source evaluation framework for pathology foundation models, for LP evaluation where available [^3] Available at https://github.com/kaiko-ai/eva, accessed on Oct 31, 2024 (version: 0.1.3). This includes the BACH aresta2019bach , CRC-100k crc-100k , MHIST mhist , and PCAM pcam datasets. Linear classification on extracted embeddings in eva is done by training a single-layer neural network with a batch size of 256 patches using stochastic gradient descent (SGD) with a cosine learning rate schedule and base learning rate of 0.0003 for a total of 12.5k iterations.
For patch-level classification tasks not implemented in eva, we use an internal LP evaluation framework. These include the MSI CRC (patch), MSI STAD (patch), Pan-cancer TIL, TCGA Uniform (10x), and TCGA Uniform (20x) datasets. In the internal framework, we perform LP on extracted embeddings using scikit-learn’s Logistic Regression scikit-learn with balanced class weights. The L2 regularization parameter is chosen by performing a cross-validated grid search over 15 different values between $10^{-8}$ and $10^{4}$. Using the best parameter, a final model is fit and applied to the test set.
@@
For all slide-level classification tasks, we use the eva kaiko.ai2024eva framework, which applies the default Attention-based Multiple Instance Learning (ABMIL) ilse2018attention protocol. Here, an ABMIL head with ReLU activations is trained with a batch size of 32 slides using the AdamW optimizer loshchilov2018decoupled with a cosine learning rate schedule and base learning rate of 0.001 for a total of 12.5k iterations. Balanced accuracy was utilized as a performance metric for all slide-level classification tasks.
@@
The HEST-Benchmark jaume2024hest is composed of tasks for gene expression prediction and designed as multivariate regression. We follow and use the default evaluation protocol as recommended and implemented [^4] Available at https://github.com/mahmoodlab/HEST, accessed git commit ’5a0cbba’. by the authors. We provide respective details in the description of the HEST-Benchmark in Appendix A.1.
@@
Group Benchmark Phikon v2 filiot_phikon-v2_2024 UNI chen_towards_2024 Gigapath xu_whole-slide_2024 RudolfV dippel_rudolfv_2024 Virchow2 zimmermann_virchow_2024 H-optimus-0 hoptimus0 Ours Molecular-related HEST-COAD 25.6 26.2 30.7 31.0 25.9 30.9 29.4 HEST-HCC 7.8 8.3 7.1 9.4 9.6 8.4 10.7 HEST-IDC 56.6 58.5 56.8 57.4 59.3 61.0 60.4 HEST-LUAD 54.8 55.2 55.8 57.7 56.9 57.3 58.0 HEST-LYMPH_IDC 24.8 25.8 25.1 25.6 25.9 26.8 26.4 HEST-PAAD 47.9 48.8 49.5 51.1 47.3 50.9 51.8 HEST-PRAD 37.7 32.2 38.4 37.7 35.1 38.5 38.4 HEST-READ 18.5 18.4 19.6 19.9 21.1 24.1 22.8 HEST-SKCM 58.4 63.5 58.8 61.8 63.7 66.1 62.5 HEST-ccRCC 27.3 25.3 24.9 25.3 27.4 29.0 29.4 MSI CRC (patch) 68.8 69.5 70.4 69.9 74.0 71.2 73.6 MSI STAD (patch) 71.2 70.5 71.0 74.1 74.8 73.6 76.0 Molecular-Average 41.6 41.8 42.3 43.4 43.4 44.8 44.9 Morphology-related Pan-cancer TIL 92.9 92.6 92.3 92.6 93.1 93.0 93.0 TCGA Uniform (10x) 64.0 68.6 69.1 70.6 73.0 70.4 71.8 TCGA Uniform (20x) 69.8 67.8 68.0 78.1 71.5 72.4 67.8 BACH 73.8 80.1 80.2 76.9 88.7 75.8 93.1 CRC-100k 95.5 95.4 95.9 96.0 96.7 96.2 97.1 MHIST 78.4 84.4 83.1 80.5 85.9 85.0 86.4 PCAM 90.0 93.6 94.5 94.6 93.9 94.3 94.9 CAMELYON16 79.8 85.0 82.1 77.1 86.5 84.0 86.8 PANDA 65.3 69.6 69.6 69.6 66.4 68.0 70.5 Morphology-Average 78.8 81.9 81.6 81.8 84.0 82.1 84.6 Overall Average 57.6 59.0 59.2 59.9 60.8 60.8 61.9
@@ 4 Results
The following analysis is based on 21 public benchmark datasets from two public foundation model evaluation frameworks eva kaiko.ai2024eva and HEST jaume2024hest as well as additional tumor-micro-environment (TME) and cancer typing benchmark datasets. The tasks range from TME tissue- and cell-typing over identifying morphological patterns, identifying cancer subtypes, to classifying molecular mutations. The task descriptions and evaluation protocols are detailed in Appendix A.1 and Section 3.3, respectively.
Our model achieves an average performance score of 61.9%, a 1.1 p.p. improvement over the two closest contenders, Virchow2 zimmermann_virchow_2024 and H-Optimus-0 hoptimus0 (see Table 1). It displayed the highest performance on 6 out of 9 morphology-related tasks with Virchow2 being the closest contender performing best on 2 tasks. For the molecular-related biomarker tasks H-Optimus and our model perform both best on 5 out of 12 tasks. Overall, our model displayed the best performance on 11 of 21 assessed tasks across both molecular- and morphology-related tasks. In 7 of the 10 benchmarks in which ours was not the leading model it was the second best model by performance. The model displayed below-average performance for a single benchmark, TCGA Uniform (20$\times$), where ours and UNI were the poorest performers; interestingly, our performance on the 10$\times$ version of that benchmark was second only to Virchow2.
@@ 5 Discussion
The presented model showed consistently good results across a diverse set of benchmarks covering molecular and morphology related tasks. Related work using a larger data base zimmermann_virchow_2024 or more model parameters zimmermann_virchow_2024 ; hoptimus0 ; xu_whole-slide_2024 suggests that scaling data and model size even further might yield additional improvements.
To accurately assess the quality, generalization, and robustness of pathology foundation models, it is essential to utilize a wide range of (public) benchmark datasets, covering the entire spectrum of diseases, morphologies, scanners, etc. Despite first standardized frameworks such as eva kaiko.ai2024eva and HEST jaume2024hest being available, foundation model development would benefit from a larger and more diverse pool of benchmark datasets. Access to such diverse datasets would provide deeper insights into model performance and robustness, thereby enhancing our understanding of the capabilities and limitations of pathology foundation models.

### v2

# Atlas: A Novel Pathology Foundation Model by Mayo Clinic, Charité, and Aignostics

@@ Abstract
Recent advances in digital pathology have demonstrated the effectiveness of foundation models across diverse applications. In this report, we present Atlas, a novel vision foundation model based on the RudolfV approach. Our model was trained on a dataset comprising $1.2$ million histopathology whole slide images, collected from two medical institutions: Mayo Clinic and Charité - Universtätsmedizin Berlin. Comprehensive evaluations show that Atlas achieves state-of-the-art performance across twenty-one public benchmark datasets, even though it is neither the largest model by parameter count nor by training dataset size.
@@ 1 Introduction
However, despite artificial intelligence (AI) having led to promising proof-of-concepts and applications (e.g., [24, 8, 33, 5, 23]), generalization, application variety, and robustness remain challenging and have hindered the broad translation of AI applications into clinical routine diagnostics. Here, the limited adoption of digitization in clinical practice results in a scarcity of training data, particularly for infrequent and rare diseases [11]. Furthermore, generating sufficient labeled data representing the full spectrum of human disease, biological, and technical variability inherent to morphology, tissue processing, staining, and slide scanners has proven logistically and financially challenging.
@@
In this report, we utilized a corpus of 1.2 million histopathology whole slide images derived from more than 490,000 cases from Mayo Clinic and Charité - Universitätsmedizin Berlin to train a ViT-H/14 pathology foundation model called “Atlas” using the training paradigm from RudolfV [10]. Atlas incorporates a broad diversity of diseases, staining types, and scanners, and utilizes multiple image magnifications during training. We provide an assessment of its performance compared to other leading models available for testing using twenty-one benchmark datasets assessing a variety of downstream pathology tasks. An overview of model characteristics and results can be found in Figure 1.
Figure 1: Overview of average performance, training dataset size, and model size of different contenders. The average performance is detailed in Table 1. H-Optimus-0 [34] and Prov-GigaPath [41] are the models with the most parameters and Virchow2 [42] is the model trained on most slides. Our model Atlas exhibits the best average performance and an intermediate model and training dataset size.
@@ 2 Related Work
Multiple previous works on histopathology foundation models have developed a variety of models with increasingly large parameter counts and on increasingly expansive pathology datasets curated from public and private sources [10, 42, 13, 41, 34, 7, 9, 26, 17, 28, 32]. Pathology foundation models can be separated into tile- and slide-based models. Tile-based models, which represent the majority of the models, generate embeddings from fixed-sized image tiles derived from gigapixel whole slide images (WSIs). Tile-based models have typically been trained via self-supervised-learning [10, 42, 13, 41, 34, 7, 1, 29, 38, 19, 28] and most studies [10, 42, 13, 41, 34, 7, 1, 29] are based on the DINOv2 framework [30]. We include the leading and available tile-based models in our study, i.e. Virchow2 (632 million parameters; 3.1 million slides; [42]), H-Optimus (1.1 billion parameters; 500k slides; [34]), RudolfV (300 million parameters; 134k slides; [10]), UNI (300 million parameters; 100k slides; [7]), and Phikon 2 (300 million parameters; 58k slides; [13]).
Slide-based foundation models seek to create global representations of whole slide images, which requires encoding features that may span individual tiles or otherwise represent global phenomena not readily evident on single tiles. To address this, most current slide-based models operate in two steps [9, 26, 39, 41, 36]. First, by encoding individual tiles and secondly by aggregating the resulting tile representations into a slide-level representation. For such models, the tile-encoder is a key performance gateway. Accordingly, this study compares the tile-based encoder of Prov-GigaPath (1.1 billion parameters; 171k slides; [41]).
To date, many foundation models in pathology have been trained exclusively on hematoxylin and eosin (H&E) stained slides with tiles extracted at a single magnification level. While H&E staining represents the bulk of routine pathology, alternative histochemical and immunohistochemical (IHC) stains are the basis for many biomarker evaluations and accordingly have been increasingly represented in recent studies [42, 10, 41]. Similarly, data from multiple image magnifications have been included in recent work [42, 1]. Atlas incorporates these features, and uses data obtained from H&E, IHC, and special stains, as well as multiple magnifications.
@@ 3.1 Dataset and Preprocessing
A curated set of 1.2 million de-identified WSIs, derived from 490k pathology cases, from the digital archives of Mayo Clinic and Charité - Universitätsmedizin Berlin, was utilized to generate 3.4 billion image tiles for training. Tiles were extracted at multiple resolutions, namely 0.25, 0.5, 1.0, and 2.0 microns per pixel, corresponding to objective microscopic magnifications of 40×40$\times$40 ×, 20×20$\times$20 ×, 10×10$\times$10 ×, and 5×5$\times$5 ×, respectively. An overview of relevant dataset statistics are given in Figure 2. Slides and tiles were sampled for training using the sampling algorithm of RudolfV [10].
@@ 3.2 Model Framework and Compute Environment
A sampled dataset of ca. 520m tiles was used to train a ViT-H/14 (632 million parameters) model [12] using an adapted RudolfV [10] approach, which is based on the DinoV2 framework [30]. Model training was performed with Nvidia H100 GPUs within the Mayo Clinic Platform environment [^1] https://www.mayoclinicplatform.org/.
@@ 3.3 Evaluation Protocols
We evaluated model performance using linear probing protocols as established in the literature [10, 34, 42], with all models evaluated on extracted embeddings from frozen backbones. We use both public benchmarks as well as public evaluation frameworks where available, to foster reproducibility and comparability. Results were computed for 5 seeds over data split, shuffling, and “probing” model initialization per foundation model and task, if not stated otherwise in a specific task description (see Appendix A.1). We report the mean performance and standard errors over the seeds. Dataset splits are described in detail in the respective task descriptions in Appendix A.1. No augmentations were applied when extracting embeddings. As all models use Vision Transformer architectures [12], we always evaluate both the CLS token and the CLS+Mean [^2] CLS+Mean being the concatenation of the CLS token and the mean over all “mini-patch” tokens token embeddings for every model and report the better (maximum) performance of the two in Table 1. This accounts for potential systematic differences between information encoded in different tokens between different models, as also done in previous works (see e.g. [42]). Results of using the CLS and CLS+Mean token only, are reported in Table 4 and Table 5, respectively.
@@
We use eva [14], an open-source evaluation framework for pathology foundation models, for LP evaluation where available [^3] Available at https://github.com/kaiko-ai/eva, accessed on Oct 31, 2024 (version: 0.1.3). This includes the BACH [3], CRC-100k [20], MHIST [40], and PCAM [37] datasets. Linear classification on extracted embeddings in eva is done by training a single-layer neural network with a batch size of 256 patches using stochastic gradient descent (SGD) with a cosine learning rate schedule and base learning rate of 0.0003 for a total of 12.5k iterations.
For patch-level classification tasks not implemented in eva, we use an internal LP evaluation framework. These include the MSI CRC (patch), MSI STAD (patch), Pan-cancer TIL, TCGA Uniform (10x), and TCGA Uniform (20x) datasets. In the internal framework, we perform LP on extracted embeddings using scikit-learn’s Logistic Regression [31] with balanced class weights. The L2 regularization parameter is chosen by performing a cross-validated grid search over 15 different values between $10^{-8}$ and $10^{4}$. Using the best parameter, a final model is fit and applied to the test set.
@@
For all slide-level classification tasks, we use the eva [14] framework, which applies the default Attention-based Multiple Instance Learning (ABMIL) [15] protocol. Here, an ABMIL head with ReLU activations is trained with a batch size of 32 slides using the AdamW optimizer [27] with a cosine learning rate schedule and base learning rate of 0.001 for a total of 12.5k iterations. Balanced accuracy was utilized as a performance metric for all slide-level classification tasks.
@@
The HEST-Benchmark [16] is composed of tasks for gene expression prediction and designed as multivariate regression. We follow and use the default evaluation protocol as recommended and implemented [^4] Available at https://github.com/mahmoodlab/HEST, accessed git commit ’5a0cbba’. by the authors. We provide respective details in the description of the HEST-Benchmark in Appendix A.1.
@@
Group Benchmark Phikon v2 [13] UNI [7] Gigapath [41] RudolfV [10] Virchow2 [42] H-optimus-0 [34] Atlas Molecular-related HEST-COAD 25.6 26.2 30.7 31.0 25.9 30.9 29.4 HEST-HCC 7.8 8.3 7.1 9.4 9.6 8.4 10.7 HEST-IDC 56.6 58.5 56.8 57.4 59.3 61.0 60.4 HEST-LUAD 54.8 55.2 55.8 57.7 56.9 57.3 58.0 HEST-LYMPH_IDC 24.8 25.8 25.1 25.6 25.9 26.8 26.4 HEST-PAAD 47.9 48.8 49.5 51.1 47.3 50.9 51.8 HEST-PRAD 37.7 32.2 38.4 37.7 35.1 38.5 38.4 HEST-READ 18.5 18.4 19.6 19.9 21.1 24.1 22.8 HEST-SKCM 58.4 63.5 58.8 61.8 63.7 66.1 62.5 HEST-ccRCC 27.3 25.3 24.9 25.3 27.4 29.0 29.4 MSI CRC (patch) 68.8 69.5 70.4 69.9 74.0 71.2 73.6 MSI STAD (patch) 71.2 70.5 71.0 74.1 74.8 73.6 76.0 Molecular-Average 41.6 41.8 42.3 43.4 43.4 44.8 44.9 Morphology-related Pan-cancer TIL 92.9 92.6 92.3 92.6 93.1 93.0 93.0 TCGA Uniform (10x) 64.0 68.6 69.1 70.6 73.0 70.4 71.8 TCGA Uniform (20x) 69.8 67.8 68.0 78.1 71.5 72.4 67.8 BACH 73.8 80.1 80.2 76.9 88.7 75.8 93.1 CRC-100k 95.5 95.4 95.9 96.0 96.7 96.2 97.1 MHIST 78.4 84.4 83.1 80.5 85.9 85.0 86.4 PCAM 90.0 93.6 94.5 94.6 93.9 94.3 94.9 CAMELYON16 79.8 85.0 82.1 77.1 86.5 84.0 86.8 PANDA 65.3 69.6 69.6 69.6 66.4 68.0 70.5 Morphology-Average 78.8 81.9 81.6 81.8 84.0 82.1 84.6 Overall Average 57.6 59.0 59.2 59.9 60.8 60.8 61.9
@@ 4 Results
The following analysis is based on 21 public benchmark datasets from two public foundation model evaluation frameworks eva [14] and HEST [16] as well as additional tumor-micro-environment (TME) and cancer typing benchmark datasets. The tasks range from TME tissue- and cell-typing over identifying morphological patterns, identifying cancer subtypes, to classifying molecular mutations. The task descriptions and evaluation protocols are detailed in Appendix A.1 and Section 3.3, respectively.
Atlas achieves an average performance score of 61.9%, a 1.1 p.p. improvement over the two closest contenders, Virchow2 [42] and H-Optimus-0 [34] (see Table 1). It displayed the highest performance on 6 out of 9 morphology-related tasks with Virchow2 being the closest contender performing best on 2 tasks. For the molecular-related biomarker tasks H-Optimus and Atlas perform both best on 5 out of 12 tasks. Overall, Atlas displayed the best performance on 11 of 21 assessed tasks across both molecular- and morphology-related tasks. In 7 of the 10 benchmarks in which Atlas was not the leading model it was the second best model by performance. The model displayed below-average performance for a single benchmark, TCGA Uniform (20$\times$), where Atlas and UNI were the poorest performers; interestingly, Atlas’ performance on the 10$\times$ version of that benchmark was second only to Virchow2.
@@ 5 Discussion
Our novel pathology foundation model Atlas showed consistently good results across a diverse set of benchmarks covering molecular and morphology related tasks. Related work using a larger data base [42] or more model parameters [42, 34, 41] suggests that scaling data and model size even further might yield additional improvements.
To accurately assess the quality, generalization, and robustness of pathology foundation models, it is essential to utilize a wide range of (public) benchmark datasets, covering the entire spectrum of diseases, morphologies, scanners, etc. Despite first standardized frameworks such as eva [14] and HEST [16] being available, foundation model development would benefit from a larger and more diverse pool of benchmark datasets. Access to such diverse datasets would provide deeper insights into model performance and robustness, thereby enhancing our understanding of the capabilities and limitations of pathology foundation models.